In a small population that is not in Hardy-weinberg equilibrium and in which the
Answer #1

The answer is as follows :

A(p2) = 52/114 = 0.45

B(p2) = 27/112 = 0.23

AB(2pq) = 35/112 = 0.30

Frequency of A : p2 + 1/2(2pq) = 0.45 +1/2(0.30) = 0.6

Frequency of B : q= 1-p = 1-0.6 = 0.4